OpenStudy (anonymous):

Solve A = 2πr + 3t for t.

OpenStudy (anonymous):

heeeeelllllllllp

jimthompson5910 (jim_thompson5910):

Subtract 2πr from both sides Then divide both sides by 3 to isolate t

jimthompson5910 (jim_thompson5910):

A = 2πr + 3t A - 2πr = 3t (A - 2πr)/3 = t \(\Large t = \frac{A - 2\pi r}{3}\)